# Supervised Score-Based Modeling by Gradient Boosting

### v1

@@ Abstract
Score-based generative models can effectively learn the distribution of data by estimating the gradient of the distribution. Due to the multi-step denoising characteristic, researchers have recently considered combining score-based generative models with the gradient boosting algorithm, a multi-step supervised learning algorithm, to solve supervised learning tasks. However, existing generative model algorithms are often limited by the stochastic nature of the models and the long inference time, impacting prediction performances. Therefore, we propose a Supervised Score-based Model (SSM), which can be viewed as a gradient boosting algorithm combining score matching. We provide a theoretical analysis of learning and sampling for SSM to balance inference time and prediction accuracy. Via the ablation experiment in selected examples, we demonstrate the outstanding performances of the proposed techniques. Additionally, we compare our model with other probabilistic models, including Natural Gradient Boosting (NGboost), Classification and Regression Diffusion Models (CARD), Diffusion Boosted Trees (DBT), and Bayesian neural network-based models. The experimental results show that our model outperforms existing models in both accuracy and inference time.
@@ Introduction
Supervised learning is another important research area in machine learning different from generative learning. In supervised learning, a model is trained on labeled data with the goal of learning a mapping from inputs to outputs, allowing it to predict the output for unseen data accurately (Cunningham, Cord, and Delany 2008). Regression and classification problems are two basic types of supervised learning tasks, which aim to train a neural network for prediction. For regression problems, this prediction result is the target output corresponding to the input data, while for classification problems, it is the accurate label. The learning performance of this data relationship is closely related to the learning ability of the sophisticated structure of neural network models (He et al. 2016). Particularly, the Gradient Boosting Machine (GBM) is a supervised learning model that combines several weak learners into a strong learner (Bentéjac, Csörgő, and Martínez-Muñoz 2021). The sequential weak learners reduce errors from previous iterations to obtain accurate predictions, especially with large and complex datasets. Recently, GBM algorithms such as XGBoost (Chen and Guestrin 2016) and LightGBM (Ke et al. 2017) have demonstrated state-of-the-art performance in tabular competitions on platforms, including Kaggle.
@@
The GBM is a supervised learning framework that combines several weak learners into strong learners in an iterative way (Friedman 2001). In GBM, a boosted model $F(x)$ is a weighted linear combination of the weak learners, which can be formulated as
$F(x)=F_{0}(x)+\sum_{k=1}^{m}\alpha_{k}h_{k}(x),$ (1)
@@
Table 2: Inference time on different ablation settings on toy examples. W/ and w/o noise represent whether the noise term is included in Langvin dynamics; fast denotes the use of Technique 2; L1 and L2 represent the degree of noise levels of 1 and 2, respectively.
@@
NGboost MC Dropout Deep Ensembles GCDS CARD DBT SSM (L1) SSM (L2)
Boston 2.94 $\pm$ 0.53 3.06 $\pm$ 0.96 3.17 $\pm$ 1.05 2.75 $\pm$ 0.58 2.61 $\pm$ 0.63 2.73 $\pm$ 0.62 2.06 $\pm$ 0.19 2.04 $\pm$ 0.14
Concrete 5.06 $\pm$ 0.61 5.09 $\pm$ 0.60 4.91 $\pm$ 0.47 5.39 $\pm$ 0.55 4.77 $\pm$ 0.46 4.56 $\pm$ 0.50 5.88 $\pm$ 1.00 3.20 $\pm$ 1.00
Energy 0.46 $\pm$ 0.06 1.70 $\pm$ 0.22 2.02 $\pm$ 0.32 0.64 $\pm$ 0.09 0.52 $\pm$ 0.07 0.52 $\pm$ 0.07 0.36 $\pm$ 0.09 0.31 $\pm$ 0.02
Kin8nm 16.03 $\pm$ 0.04 7.10 $\pm$ 0.26 8.65 $\pm$ 0.47 8.88 $\pm$ 0.42 6.32 $\pm$ 0.18 7.04 $\pm$ 0.23 8.30 $\pm$ 0.10 4.66 $\pm$ 0.30
Naval 0.13 $\pm$ 0.00 0.08 $\pm$ 0.03 0.09 $\pm$ 0.01 0.14 $\pm$ 0.05 0.02 $\pm$ 0.00 0.07 $\pm$ 0.01 0.10 $\pm$ 0.00 0.01 $\pm$ 0.00
Power 3.79 $\pm$ 0.18 4.04 $\pm$ 0.14 4.02 $\pm$ 0.15 4.11 $\pm$ 0.16 3.93 $\pm$ 0.17 3.95 $\pm$ 0.16 4.25 $\pm$ 0.01 3.67 $\pm$ 0.33
Protein 4.33 $\pm$ 0.03 4.16 $\pm$ 0.12 4.45 $\pm$ 0.02 4.50 $\pm$ 0.02 3.73 $\pm$ 0.01 3.81 $\pm$ 0.04 5.20 $\pm$ 0.01 3.95 $\pm$ 0.43
Wine 0.63 $\pm$ 0.04 0.62 $\pm$ 0.04 0.63 $\pm$ 0.04 0.66 $\pm$ 0.04 0.63 $\pm$ 0.04 0.61$\pm$ 0.04 0.68 $\pm$ 0.01 0.46 $\pm$ 0.21
Yacht 0.50 $\pm$ 0.20 0.84 $\pm$ 0.27 1.19 $\pm$ 0.49 0.79 $\pm$ 0.26 0.65 $\pm$ 0.25 1.08 $\pm$ 0.39 0.56 $\pm$ 0.16 0.37 $\pm$ 0.21
Year 8.94 $\pm$ NA 8.77 $\pm$ NA 8.79 $\pm$ NA 9.20 $\pm$ NA 8.70 $\pm$ NA 8.81 $\pm$ NA 10.14 $\pm$ NA 10.04 $\pm$ NA
# best 0 0 0 0 2 0 0 8

### v2

@@ Abstract
Score-based generative models can effectively learn the distribution of data by estimating the gradient of the distribution. Due to the multi-step denoising characteristic, researchers have recently considered combining score-based generative models with the gradient boosting algorithm, a multi-step supervised learning algorithm, to solve supervised learning tasks. However, existing generative model algorithms are often limited by the stochastic nature of the models and the long inference time, impacting prediction performances. Therefore, we propose a Supervised Score-based Model (SSM), which can be viewed as a gradient boosting algorithm combining score matching. We provide a theoretical analysis of learning and sampling for SSM to balance inference time and prediction accuracy. Via the ablation experiment in selected examples, we demonstrate the outstanding performances of the proposed techniques. Additionally, we compare our model with other probabilistic models, including Natural Gradient Boosting (NGboost), Classification and Regression Diffusion Models (CARD), Diffusion Boosted Trees (DBT), and non-probabilistic GBM models. The experimental results show that our model outperforms existing models in both accuracy and inference time.
@@ Introduction
Supervised learning is another important research area in machine learning different from generative learning. In supervised learning, a model is trained on labeled data with the goal of learning a mapping from inputs to outputs, allowing it to predict the output for unseen data accurately (Cunningham, Cord, and Delany 2008). Regression and classification problems are two basic types of supervised learning tasks, which aim to train a model for prediction. For regression problems, this prediction result is the target output corresponding to the input data, while for classification problems, it is the accurate label. The learning performance of this data relationship is closely related to the learning ability of the sophisticated structure of neural network models (He et al. 2016). Particularly, the Gradient Boosting Machine (GBM) is a supervised learning model that combines several weak learners into a strong learner (Bentéjac, Csörgő, and Martínez-Muñoz 2021). The sequential weak learners reduce errors from previous iterations to obtain accurate predictions, especially with large and complex datasets. Recently, GBM algorithms such as XGBoost (Chen and Guestrin 2016) and LightGBM (Ke et al. 2017) have demonstrated state-of-the-art performance in tabular competitions on platforms, including Kaggle.
@@
The GBM is a supervised learning framework that combines several weak learners into strong learners in an iterative way (Friedman 2001). In GBM, a boosted model $F(x)$ is a weighted linear combination of $m$ weak learners, which can be formulated as
$F(x)=F_{m}(x)=F_{0}(x)+\sum_{k=1}^{m}\alpha_{k}h_{k}(x),$ (1)
@@
Experiment
Toy examples
To demonstrate the effectiveness of SSM and all improved techniques, we first perform experiments on 5 selected toy examples (linear regression, quadratic regression, log-log linear regression, log-log cubic regression, and sinusoidal regression) proposed in (Han, Zheng, and Zhou 2022). We add unbiased normal distribution noise to the original models to obtain noisy data for training, increasing model complexity. We aim to predict accurate regression results for pure input data in the inference process. To evaluate the impact of various technologies, we have designed an ablation experiment that considers three variables: the use of original Langevin dynamics, the implementation of the fast sampling technique as described in Technique 2, and the application of different coefficients $\lambda(\sigma_{i})$ in the loss function, where L1 and L2 represent the degree of noise levels of 1 and 2, respectively. The experiment results with the Root Mean Squared Error (RMSE) and inference time about different model settings are shown in Table 2 and 3, respectively. The scatter plots for 5 examples of both true and generated data following the first setting, i.e., L1, w/o noise fast, are shown in Figure 1. We observe that while perturbed data is used, SSM can still fit the regression equation of the original data. The proposed fast sampling Technique 2 significantly reduces the inference time for all tasks. Additionally, since the fast sampling technique ensures the error bound of the current estimated state, it can even achieve the best performance on tasks except for log-log linear regression. Using the noise-free Langevin equation, i.e., w/o noise, performs better on most tasks, indicating that the selection of different coefficients also affects the performance of the model.
Regression
We further evaluate our model on 10 UCI regression tasks (Dua and Graff 2017). We employ the same experimental settings as those used in the CARD model (Han, Zheng, and Zhou 2022). We compare our model with NGboost (Duan et al. 2020), CARD (Han, Zheng, and Zhou 2022), DBT (Han and Zhou 2024), non-probabilistic GBM models, including CatBoost (Prokhorenkova et al. 2018) and XGBoost (Chen and Guestrin 2016), and GCDS (Zhou et al. 2023). The experiment results with RMSE are shown in Table 4. It is noticed that SSM outperforms existing approaches in RMSE. SSM (L2) obtains the best results in 8 out of 10 datasets. Although the RMSE of SSM (L1) is larger, it has a smaller standard deviation, which means the network is more stable. In addition, probabilistic models such as CARD predict the probability of distribution, while SSM predicts the points directly. This also leads to a lower standard deviation of SSM on most tasks.
Classification
Model CARD SSM (L1) SSM (L2)
CIFAR Accuracy 90.93 $\pm$ 0.02 90.91 $\pm$ 0.00 90.99 $\pm$ 0.00
-10 Time 50.30 $\pm$ 0.30 19.63 $\pm$ 1.21 15.43 $\pm$ 1.12
CIFAR Accuracy 71.42 $\pm$ 0.01 71.51 $\pm$ 0.00 71.38 $\pm$ 0.00
-100 Time 90.58 $\pm$ 0.12 32.30 $\pm$ 0.60 27.16 $\pm$ 0.19
Table 1: Accuracy (in %) and inference time (in second (s)) on CIFAR-10 and CIFAR-100.
@@
Table 3: Inference time on different ablation settings on toy examples. W/ and w/o noise represent whether the noise term is included in Langevin dynamics; fast denotes the use of Technique 2; L1 and L2 represent the degree of noise levels of 1 and 2, respectively.
@@
NGboost CatBoost XGBoost GCDS CARD DBT SSM (L1) SSM (L2)
Boston 2.94 $\pm$ 0.53 2.19 $\pm$ 0.09 2.23 $\pm$ 0.04 2.75 $\pm$ 0.58 2.61 $\pm$ 0.63 2.73 $\pm$ 0.62 2.06 $\pm$ 0.19 2.04 $\pm$ 0.14
Concrete 5.06 $\pm$ 0.61 4.75 $\pm$ 0.16 4.62 $\pm$ 0.13 5.39 $\pm$ 0.55 4.77 $\pm$ 0.46 4.56 $\pm$ 0.50 5.88 $\pm$ 1.00 3.20 $\pm$ 1.00
Energy 0.46 $\pm$ 0.06 0.32 $\pm$ 0.02 0.23 $\pm$ 0.01 0.64 $\pm$ 0.09 0.52 $\pm$ 0.07 0.52 $\pm$ 0.07 0.36 $\pm$ 0.09 0.31 $\pm$ 0.02
Kin8nm 16.03 $\pm$ 0.04 9.31 $\pm$ 0.15 11.89 $\pm$ 0.17 8.88 $\pm$ 0.42 6.32 $\pm$ 0.18 7.04 $\pm$ 0.23 8.30 $\pm$ 0.10 4.66 $\pm$ 0.30
Naval 0.13 $\pm$ 0.00 0.13 $\pm$ 0.00 0.10 $\pm$ 0.00 0.14 $\pm$ 0.05 0.02 $\pm$ 0.00 0.07 $\pm$ 0.01 0.10 $\pm$ 0.00 0.01 $\pm$ 0.00
Power 3.79 $\pm$ 0.18 3.41 $\pm$ 0.03 3.22 $\pm$ 0.03 4.11 $\pm$ 0.16 3.93 $\pm$ 0.17 3.95 $\pm$ 0.16 4.25 $\pm$ 0.01 3.67 $\pm$ 0.33
Protein 4.33 $\pm$ 0.03 3.83 $\pm$ 0.02 3.53 $\pm$ 0.01 4.50 $\pm$ 0.02 3.73 $\pm$ 0.01 3.81 $\pm$ 0.04 5.20 $\pm$ 0.01 3.95 $\pm$ 0.43
Wine 0.63 $\pm$ 0.04 0.58 $\pm$ 0.02 0.57 $\pm$ 0.01 0.66 $\pm$ 0.04 0.63 $\pm$ 0.04 0.61$\pm$ 0.04 0.68 $\pm$ 0.01 0.46 $\pm$ 0.21
Yacht 0.50 $\pm$ 0.20 0.44$\pm$ 0.07 0.43 $\pm$ 0.08 0.79 $\pm$ 0.26 0.65 $\pm$ 0.25 1.08 $\pm$ 0.39 0.56 $\pm$ 0.16 0.37 $\pm$ 0.21
Year 8.94 $\pm$ NA 8.94 $\pm$ NA 8.77 $\pm$ NA 9.20 $\pm$ NA 8.70 $\pm$ NA 8.81 $\pm$ NA 10.14 $\pm$ NA 10.04 $\pm$ NA
# best 0 0 3 0 1 0 0 6